Please help me with following. How do I set it up?
Sample has 4oz of A and 3 oz of B material. Have to add C to this mixture to make a final mixture that is 50%A and 50%B. C is composed of 33%A and 66%B. How much C (oz)have to be added?

Answer by stanbon(75887)

Sample has 4oz of A and 3 oz of B material. Have to add C to this mixture to make a final mixture that is 50%A and 50%B. C is composed of 33%A and 66%B. How much C (oz)have to be added?
-------
Let x be athe amount of C added to the original mix.
(1/3) of x is A
(2/3) of x is B
----
To get a 50-50 mix
4 + (1/3)x = 3+(2/3)x
---
(1/3)x = 1
x = 3 oz
------
Ans: Amt. of C to add is 3 ounces
------------------------------------
Cheers,
Stan H.
======================
Cheers,
Stan H.
----------------